OpenStudy (anonymous):

Determine 0. sqrt3 tan0 = 1

OpenStudy (anonymous):

Sorry by 0 I mean degree/ So it is telling me to find the degree. the answer is 30 degrees, but I dont know how

OpenStudy (anonymous):

\[\sqrt{3}\tan \theta =1\] \[\tan \theta =1/\sqrt{3}\] so... \[\theta = \tan ^{-1}1/ \sqrt{3}\] plug it into your calculator and you get \[\theta = 30\]

OpenStudy (anonymous):

wow thank you so much ! :D